OpenStudy (swissgirl):

Reduce the polynomial and find its complex roots \(f(x)=x^4+5x^3-9x^2-85x-136\)

OpenStudy (swissgirl):

lol didnt know u were on

OpenStudy (swissgirl):

not quite sure how to reduce it since u start dealing with decimals. It shld be quite simple though

OpenStudy (kinggeorge):

By reduce do you mean factor?

OpenStudy (swissgirl):

i am assuming so

OpenStudy (kinggeorge):

Well, I have a couple ideas, so let's see if I can do this by hand... \[f(x)=x^4+5x^3-9x^2-85x-136\\ f(x)=x^4-17x^2+5x^3+8x^2-17(5x-8)\\ f(x)=x^4-17x^2+x^2(5x+8)-17(5x+8)\\ f(x)=x^4-17x^2+(x^2-17)(5x+8)\\ f(x)=x^2(x^2-17)+(x^2-17)(5x+8)\\ f(x)=(x^2-17)(x^2+5x+8) \]

OpenStudy (kinggeorge):

So you can use the quadratic formula on \(x^2+5x+8\) to find the complex roots, and just \[x^2-17=0\implies x=\pm\sqrt{17}\]For the real roots.

OpenStudy (swissgirl):

Let me just see if I can follow what u did

OpenStudy (kinggeorge):

Take your time. I'm honestly very surprised I was able to factor that almost entirely by hand. (I used wolfram to check correctness, and to find the gcd of 85, 136).

OpenStudy (swissgirl):

idk if its possible. Dont drive urself crazy buttt I cant decipher what is squred and what is tripled. Is it possible to make it larger

OpenStudy (swissgirl):

whoops cubed lol

OpenStudy (kinggeorge):

\[\Large f(x)=x^4+5x^3-9x^2-85x-136\\ \Large f(x)=x^4-17x^2+5x^3+8x^2-17(5x-8)\\ \Large f(x)=x^4-17x^2+x^2(5x+8)-17(5x+8)\\ \Large f(x)=x^4-17x^2+(x^2-17)(5x+8)\\ \Large f(x)=x^2(x^2-17)+(x^2-17)(5x+8)\\ \Large f(x)=(x^2-17)(x^2+5x+8)\]

OpenStudy (swissgirl):

Thanks :)

OpenStudy (kinggeorge):

welcome

OpenStudy (swissgirl):

OKKKK so like will this method always work?

OpenStudy (kinggeorge):

Not always. This is just factoring by grouping, and while potentially useful, it is often difficult to identify the correct time to use it, and what to group together. I just got lucky that I guessed the right things to group. For example, if I had \[\Large f(x)=x^4+5x^3-10x^2-85x-136\]instead, I would not be able to group it as easily.

OpenStudy (swissgirl):

ohhhh i seee. I guesss I have to play around. Not as simple as I wish it wld be. Thanks KG:)))))))))))

OpenStudy (kinggeorge):

You're welcome. I will note however, that if they tell you to factor something, it's usually possible to do it by hand with a little cleverness.

OpenStudy (swissgirl):

Haha So I hope and pray that it will be so. lol Thanks KGGGG